10/29/2013 Minutes


Dominick, Kanika, Mat, Gavin, Nick, Susan, Michael


Kanika was on shift over the weekend and is now on swing shift this weekend.
Steve collects the list of FEBs that do not receive hits and passes info onto Rick that
gets discussed at Operations meetings, i.e. do we replace them.
FarDet has been running stable with cooled, unmangled, no DCMs
End of shift form has been filled in by all shifters - a great thing!
Encourage shifters to check each other.


Dominick now monitors production status. Run 11519 finished yesterday but files are not appearing.
Files are now appearing again. FTS issues for the last 24 hours. What exactly??

*Action Item: Requested to add number of raw files processed. Dominick will do this
*Action Item: Gavin to find out why FTS was down for 24 hours